OpenStudy (anonymous):

show that matrix A and transpose of A have same eigenvalues. Give an example 2by2 matrix A, where A and A^T have different eigenvectors

OpenStudy (anonymous):

\[A= \left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] \[A^T=\left[\begin{matrix}a & c \\ b & d\end{matrix}\right]\] these two matrices have same eigenvalues but different eigenvectors